OpenStudy (anonymous):

How many solutions does this system have? -3x + 6y=10 -3x + 6y=-4 A. none B. one C. infinite D. two

OpenStudy (zzr0ck3r):

since the left hand sides of the equations are exactly the same, you will always get the same number when you plug in for x and y, so it is impossible for them to equal different numbers.

OpenStudy (anonymous):

is it one?

OpenStudy (zzr0ck3r):

if they cant ever equal different things, and we have them set to equal different things, then there can be no solutions

OpenStudy (zzr0ck3r):

multiply the first equation by -1 we now have \(3x-6y=-10\\-3x+6y=-4\) Now add them together \(-3x+3x-6y+6y=-14\\0=-14\\\ \\\huge{\text{YIKES!}}\)

OpenStudy (zzr0ck3r):

multiply the first equation by -1 we now have \(3x-6y=-10\\-3x+6y=-4\) Now add them together \(-3x+3x-6y+6y=-14\\0=-14\\\ \\\huge{\text{YIKES!}}\)